OpenStudy (anonymous):

In the equation x + kx+ 54 one root is equal to twice the other root. The value(s) of k is (are): a) 5.2 15.6 -5.2 15.6 22.0

OpenStudy (caozeyuan):

Well, x+kx+54 is polynomial, not equation, so where is the equation?

OpenStudy (anonymous):

x + kx+ 54=0

OpenStudy (caozeyuan):

Cool, now let's do this

OpenStudy (caozeyuan):

the first thing I would do is solve for x as a function of k

OpenStudy (caozeyuan):

Wait a sec! this is first order, meaning it's got only one root. are you missing a square?

OpenStudy (anonymous):

oh sry, x^2

OpenStudy (caozeyuan):

So we have 2 eqa now.\[x ^{2}+kx+54=0\]

OpenStudy (caozeyuan):

and \[4x ^{2}+2kx+54=0\]

OpenStudy (caozeyuan):

with me so far?

OpenStudy (anonymous):

why are there 2 equations

OpenStudy (caozeyuan):

Becuase it says one rt is twice the other, so if one rt is x, the other has to be 2x, plug 2x in into the original eqa, we can get a new eqa that is equally valid

OpenStudy (anonymous):

oh ok

OpenStudy (caozeyuan):

now divide the new eqa by 4 on both side, give me the result

OpenStudy (anonymous):

x^2+kx/2 +27/2

OpenStudy (caozeyuan):

fantastic, subtract this eqa above from the original one, what do you get?

OpenStudy (anonymous):

why?

OpenStudy (caozeyuan):

Because you can get rid of the annoying x squared now

OpenStudy (anonymous):

@caozeyuan sorry, i had to do something. so now i set them equal to eachother and solve for x?

OpenStudy (caozeyuan):

NONONO!

OpenStudy (caozeyuan):

you subtract!

OpenStudy (radar):

The roots are:\[-3\sqrt{3}, -6\sqrt{3}\] . This means:\[k=9\sqrt{3}\] Now to verify that: \[(x + 6 \sqrt{3}) (x + 3\sqrt{3}) = x ^{2}+ 9\sqrt{3}+54\] \[(x +6\sqrt{3})=0\]\[x=-6\sqrt{3}\]\[x = -3\sqrt{3}\]

OpenStudy (caozeyuan):

@radar, show steps don't just give answers

OpenStudy (anonymous):

wait, shouldn't the second equation be 2x^2+2kx+54?

OpenStudy (caozeyuan):

well, no. if x is 2x, then x^2 is (2x)^2 which is 4x^2

OpenStudy (anonymous):

oh right, so i get kx/2 +24 after subtracting

OpenStudy (anonymous):

=0

OpenStudy (anonymous):

then what??

OpenStudy (caozeyuan):

I got kx+81=0, WTF is going on?!

OpenStudy (anonymous):

oh i forgot it's 27/2

OpenStudy (anonymous):

then what

OpenStudy (caozeyuan):

then it sucks, cuz our answer is not in the choices

OpenStudy (anonymous):

..................

OpenStudy (anonymous):

@phi @Greg_D @mathmate

OpenStudy (caozeyuan):

My brain is rusty, I haven't had any math class for half an year

OpenStudy (anonymous):

the answer found by @radar is \(9\sqrt{3}\approx 15.59\) which is close to \(15.6\), the second choice

OpenStudy (caozeyuan):

But what's wrong w/ my way?! IDC anyway cuz I finished highschool algebra a long time ago

OpenStudy (anonymous):

Howw?

OpenStudy (anonymous):

@caozeyuan please give me a minute to read the entire post...

OpenStudy (anonymous):

i dont see any problems starting with \(P(x)=0\) and \(P(2x)=0\), but when you substract you may be missing a solution, because the \(x^2\) term dissapears

OpenStudy (caozeyuan):

I see, so I guess I have to solve the eqa with quadratic formula then find k?

OpenStudy (anonymous):

let me try something in paper, if it works i will post it... just 2 minutes please

OpenStudy (anonymous):

Maybe we have to use sum and product rules? C/a and -b/a

OpenStudy (anonymous):

@caozeyuan exactly that. if we solve directly we get: \[x_{\pm} = \frac{-k\pm \sqrt{k^2-216}}{2}\] so we can try either: \(x_+ = 2x_-\) or \(x_- = 2x_+\)

OpenStudy (caozeyuan):

yep, that's my gut feeling actually, but I though subtraction is much easier

OpenStudy (anonymous):

@caozeyuan when dealing with quadratic eqs. (or higher order) you have to be careful when you substract and/or divide, you may loose solutions!! now, if we go with: \(x_+=2x_-\) we get: \[\frac{-k+\sqrt{k^2-216}}{2}=-k-\sqrt{k^2-216}\] so \[k=3\sqrt{k^2-216}\] or \[k^2=9(k^2-216)\] which solves to: \(k=-9\sqrt{3}\approx -15.58\) and \(k=9\sqrt{3}\approx 15.58\) so i guess the answer is the second choice \(k=15.6\)

OpenStudy (caozeyuan):

cool, I got that, but I don't care LOL

OpenStudy (anonymous):

lol

OpenStudy (radar):

Sorry about that. I was using the factor method to solve your quadratic equation. The two factors would be of the form (x +a)(x+b) The two factors when multiplied would have to equal your equation\[x ^{2}+kx+54=0\]The product of the numerical terms would have to equal 54. And as stated by the problem b would equal 2b ("one root is equal to twice the other root" and a + 2b would have to equal k.